OpenStudy (anonymous):

g(x)=(square root) 6-2x find the domain

OpenStudy (anonymous):

domain is all your x's, therefore solve for x: 6-2x=0 6=2x 6/2=x x=3

OpenStudy (anonymous):

sq rt(6-2x) is valid only when 6-2x>=0 implies 2x<=6 implies x<=3 hence domain (-infinity,3]

OpenStudy (anonymous):

@ArtezJ

OpenStudy (anonymous):

ohh yes missed a very important step , thanks princer_jones